# MineLand: Simulating Large-Scale Multi-Agent Interactions with Limited Multimodal Senses and Physical Needs

### v1

@@ 3.1 Programmatic Task
This type of task has a clear success criterion so that MineLand directly determines whether the task is completed based on the environmental feedback. We referred to MineDojo for the design of programmatic tasks. In total, MineLand has 4499 programmatic tasks.
Each Task $T$ is defined as a 5-tuple: $T=(G,\mathcal{G},\mathcal{I},f_{suc},S)$. $G$ refers to the task goal that needs to be completed, which will be presented in the form of a short sentence, for example: “Harvest 1 cobblestone with 1 wooden pickaxe”. $\mathcal{G}$ is guidance, which refers to the steps taken to achieve the task goal. These steps are automatically generated by OpenAI’s gpt-4-1106-preview based on the goal. It is presented in the form of several sentences. $\mathcal{I}$ is the initial condition of the task, for example, initial inventory, spawn position, and nearby enemies. $f_{suc}$ is the Success Criterion used to determine whether a task is successful. $S$ is a set of parameters set by developers to adjust task settings. These parameters include the number of agents, cooperative mode, competitive mode, etc.
@@ 3.2 Creative Task
Compared with programmatic tasks, there is no clear success criterion for creative tasks. Creative tasks is defined by a 4-tuple: $T=(G,\mathcal{G},\mathcal{I},S)$. MineLand’s creative tasks include all of the tasks in MineDojo. Besides, we construct 12 new tasks. There are 1536 creative tasks in total.
@@ 3.3 Hybrid Task
Hybrid task combines the characteristics of programmatic tasks and creative tasks. These tasks do not have a unique ground truth but have some references. We represent the hybrid task as $T=(G,\mathcal{G},\mathcal{I},\mathcal{D},f_{score},S)$. where $\mathcal{D}$ denotes the references. Unlike programmatic tasks, because Hybrid tasks do not have a ground truth, MineLand will return a score of $f_{score}$ based on $\mathcal{D}$. The higher the score, the better the task is completed. $\mathcal{D}$ can be some key rules, constraints, or key evaluation indicators.
Hybrid Tasks have two types of tasks: Construction Tasks and Stage Performance Tasks. Construction Task aims to provide a blueprint for a building or scene, and agents are expected to build these buildings or scenes based on the blueprint. The blueprint is the reference $\mathcal{D}$, and MineLand gives a score based on whether the constructed buildings meet the blueprint’s expectations. Blueprints can be pictures of buildings and scenes in real life, or Minecraft-style pictures. Moreover, as blueprints may be partial or 2D, it is meaningless to directly compare 3D buildings with 2D drawings. We use ORB (Rublee et al., 2011) to calculate the similarity between the pictures of the constructed buildings and the pictures of the blueprint. ORB $f_{score}$ is calculated by the formula of $f_{score}\in[0,1]$ represents the similarity between the constructed building picture and the blueprint.
Stage Performance Tasks aim to provide a script of a drama consisting of several behaviors, which may be a single action or an emotional expression. Agents need to perform the script. MineLand provides a comprehensive score for agents based on the accuracy of the Agent’s actions in terms of time, order, and correlation. The metrics of Stage Performance Tasks are as follows. $Seq_{agent}$ represents the action sequence generated by the agent, while $Seq*$ is the ground truth. Let $n$ and $m$ be the lengths of $Seq*$ and $Seq_{agent}$ respectively. And let $LCS$ be the length of the Longest Common Subsequence of $Seq*$ and $Seq_{agent}$. After that, the final score is calculated as $f_{score}=\frac{LCS}{n}+\frac{LCS}{m}$.
@@ 4 Alex Agent
Crafting an AI agent for MineLand, where daily life seamlessly blends with task-oriented activities, opens up exciting possibilities. The agent should fulfill daily needs like cooking, socializing, and maintaining shelter, while also completing assigned tasks like resource gathering, crafting, or construction. We propose Alex [^6] Alex is the protagonist in the sandbox game Minecraft, one of the default skins for players and a character in the game: https://www.minecraft.net/zh-hans. To pay tribute, we named our proposed AI agent Alex., a VLM-based approach as shown in Figure 4, to balance daily routines and tasks. Alex supports both individual daily-life goals and community-based task-oriented goals.
Conventional LLM-powered AI agents depend on LLM to operate as its brain, which is backed by several vital components that perform various essential functions. These components, such as the memory component, planning component, and acting component, have been thoroughly studied recently. To cater to our specific requirements, we have improved these three components (refer to Appendix C for more details) and introduced one new component: the multi-tasking component. Additionally, Alex exhibits different personality traits predefined in the system prompt.
Figure 4: Illustration of the architecture of Alex.
@@
During multitasking, people often switch attention between tasks, for example cooking while talking. The ability to communicate smoothly with other players while working on a task-oriented action is crucial in multi-agent scenarios. Therefore, we develop mechanism of multi-tasking ability to enhance the agent’s attention control and working memory abilities inspired by Multitasking theory from Cognition field (Salvucci & Taatgen, 2008). With the multi-tasking mechanism, Alex allows for simultaneously simulating and executing intricate coordination and scheduling with multiple tasks.
For attention control, Alex effectively controls attention among multiple tasks. For working memory, Alex maintains and processes information in the Memory Library. When another agent says hello to the agent, this involves saving and restoring internal states when switching between communication activities and goal-driven actions to avoid disrupting progress or forgetting ongoing tasks.
@@ 5.1 Experiments of Simulators Performance
We evaluate the number of agents that MineLandcan support and compare MineLandwith other popular Minecraft simulators. We utilize a mainstream consumer desktop PC equipped with an Intel i5-12400F CPU and 64GB of memory. Performance Monitor is employed to monitor the process. Our findings reveal that MineLandis capable of supporting up to 16 agents simultaneously while providing a visual display. When visual display is disabled, the number of concurrently running agents increases to 48. Furthermore, as depicted in Table 1, when MineLand and Malmo both run 8 agents, MineLand’s CPU and memory usage are approximately 1/3 that of Malmo’s (specifically, 35.6% and 38.0%, respectively). It is worth noting that Malmo serves as the foundation for most popular Minecraft Platforms (e.g., MineDojo/MineRL/MarLÖ(Perez-Liebana et al., 2019)), thus highlighting MineLand’s superior performance compared to the vast majority of existing Minecraft Platforms. Consequently, MineLandproves to be highly suitable for multi-agent environments.
Table 1: Part of comparison table of performance of Minecraft simulators. The full table is in Table A.1. MineLandheadless means MineLand without vision. CPU and MEM both means the average of CPU time and memory usage during initialization phase and five-minute run.
@@
5.2 Experiments of Multimodal Observation
We leverage OpenAI’s gpt-4-vision-preview API for text completion and text-embedding-ada-002 API for text embedding. The temperature is set to 0. All experiments. Unless specified, all experiments in Section 5 are set to this default setting. See Appendix D for details.
[…]
simulator time success rate
MineLand 46.38s 80%
[…]
@@ 5.3 Experiments of Multitasking
To validate the impact of multitasking support in the simulator, we conduct the obsidian mining task, which takes over 8 minutes and requires multiple steps to complete. During these steps, two types of events activate the multi-tasking component and interrupt the mining process:
Chat event: Another agent nearby initiates a conversation with Alex. This is a low-priority event, and Alex can choose whether to respond to the other agent.
Hurt event: The agent gets hurts. For example, a zombie attacks the agent. This is a high-priority event, requiring Alex to stop its current task and address this event first.
@@
5.4 Experiments of Limited Senses
To assess the impact of limited senses, we carry out the “shearing the sheep” task. An agent is required to locate a sheep to complete the task. In the case of an agent with limited visual senses, if there are no sheep within its field of view, it cannot shear a sheep. Therefore, the agent communicates with other agents to find the whereabouts of a sheep. On the contrary, an agent with unlimited visual senses is cognizant of the presence of a sheep in another location, prompting it to explore the surroundings to locate one.
5.5 Experiments of Physical Needs
We perform experiments using a survival task. The starting condition of the agent is a health level of 1 and a satiety level of 0, with an initial inventory of 2 breads. The objective for the agent is to survive for one day. Comparative experiments are carried out in regular MineLand and its counterpart version MineLand ${}_{w/o\ physical\ needs}$ without physical needs. Alex in MineLand can survive one day in his shelter and successfully complete the task while the same agent in MineLand ${}_{w/o\ physical\ needs}$ failed to execute the task. The failure is due to the lack of physical needs, although Alex is aware of possessing bread, the lack of physical needs resulted in it not consuming the bread. Instead, the agent continuously collected resources to cope with threats. Due to the low health level, it was killed by a zombie at night, leading to the failure of the task. We may refer to Appendix G for more details of short-term plans and actions generated chronologically in MineLand ${}_{w/o\ physical\ needs}$ and MineLand.
Repeated experiments on the constructed building task also show that Alex in MineLand with physical needs live longer. Agents tend to prioritize eating and then building a shelter, while Alex in MineLand without physical needs lean towards exploring and collecting resources. This indicates that physical needs’ importance to simulate real life.
5.6 Experiments of Multi-Agent Cooperation
5.7 Experiments of Single Agent
We assess individual Alex’s capabilities based on tech tree tasks, demonstrating that Alex’s architecture can plan and execute complex tasks. We attempt the task of obtaining diamonds six times, unlocking crucial items such as the crafting table, wooden pickaxe, stone pickaxe, iron ore, coal, furnace, iron ingot, and iron pickaxe in the process. Alex get diamonds twice out of six tasks. Importantly, to approach the real world, we’ve added a restriction on multimodal information atop Voyager, where Alex can “cheat” (i.e., obtain the location of the target) only when it can visually discover or reason about the target’s presence nearby.
As shown in Appendix F, even with this restriction, Alex demonstrates a strong ability for long-term planning. Additionally, when a plan does not yield the desired results (e.g., suddenly encountering obstacles on the road), Alex promptly adjusts and devises a short-term plan using its multimodal information.
5.8 Social Dynamics
In the “unlocking tools” task with two agents, two agents work together effectively, with a reduced workload per agent and higher communication expenses. On the other hand, two competitive agents worked independently and necessitated fewer code iterations. The primary reason for this is that, in adversarial relationships, agents tend to achieve more in a single iteration to expedite progress and outperform their opponents. However, this results in less thorough planning and more code errors. Consequently, multiple agents in adversarial relationships require fewer code iterations but make more mistakes.
We also observe that personality plays a significant role in determining the behavior of agents in multi-agent societies. We assigned the personality trait of high openness to both agents. Under this condition, the agents tended to establish collaboration and engage in mutual communication. When no personality was set for the agents, they tended to work independently.
Table 4: The number of code iterations needed per agent to unlock tools made of various materials is determined under three conditions. These conditions include a single agent, two agents in a cooperative relationship, and a competitive relationship. Each experiment is repeated three times, and the success rate is 100%.
Relationship Wooden Stone Iron
Single Agent 7±2 10±3 25±7
Cooperative 13±5 20±7 49±10
Competitive 6±2 10±3 27±10
@@ 6 Conclusion
Traditional multi-agent simulators have facilitated rich explorations of collective behavior, often under the assumption of perfect information and limitless capabilities. We propose a multi-agent simulator that bridges this gap through three aspects: large-scale agents, limited multimodal senses, and physical needs. The advantage carries potential broader impacts across various domains as discussed in Appendix A.
Appendix A Broader Impact

### v2

@@ 3.1 Programmatic Task
We follow MineDojo [16] for the design of programmatic tasks. Each Task $T$ is defined as a 5-tuple: $T=(G,\mathcal{G},\mathcal{I},f_{suc},S)$. $G$ refers to the task goal that needs to be completed. $\mathcal{G}$ is guidance. $\mathcal{I}$ is the initial condition of the task. $f_{suc}$ is the Success Criterion. $S$ is a set of parameters that could be customized. Different from MineDojo, these parameters include the number of agents, cooperative mode, competitive mode, etc. In total, MineLand has 4499 programmatic tasks.
@@ 3.2 Creative Task
Creative tasks is defined by a 4-tuple: $T=(G,\mathcal{G},\mathcal{I},S)$. There are 1536 creative tasks in total, and is compatible with tasks in MineDojo.
@@ 3.3 Hybrid Task
Hybrid tasks do not have a unique ground truth but have several references [^5] Hybrid tasks resemble translation tasks in that, while a single unique translation may not exist, multiple reference translations can guide the process. Here references could be key rules, constraints, and key evaluation indicators, etc.. We represent the hybrid task as $T=(G,\mathcal{G},\mathcal{I},\mathcal{D},f_{score},S)$. where $\mathcal{D}$ denotes the references. Unlike programmatic tasks, because Hybrid tasks do not have a ground truth, MineLand will return a score of $f_{score}$ based on $\mathcal{D}$. The higher the score, the better the task is completed. We design two types of tasks: Construction and Stage Performance, for hybrid tasks.
Construction Tasks Given a blueprint for a building or scene, the Construction Task aims to build these buildings or scenes based on the blueprint. The blueprint, either pictures in real life or Minecraft-style pictures, is the reference $\mathcal{D}$.
Evaluation Metrics. MineLand gives a score based on whether the constructed buildings meet the blueprint’s expectations. We calculate the task scores through the VLM-based evaluations and human evaluation, which both use the same criteria (refer to Section Q for more details). The evaluation scores range from 1 to 5, with higher scores indicating that the agents’ constructions more closely resemble the intended blueprint.
Stage Performance Tasks Given a script of a drama consisting of several behaviors, which may be a single action or an emotional expression, the Stage Performance Task aims to perform the script with agents as actors.
Evaluation Metrics. We leverage two complementary evaluation metrics: an LCS (Longest Common Subsequence)-based metric and human evaluation. The LCS-based metric consists of two distinct scores: the keypoint score and the appropriateness score. We present the formulas for adding up these two scores:
$f_{key}+f_{appro}=\frac{|LCS(SEQ_{Agent},SEQ^{*})|}{|SEQ^{*}|}+\frac{|LCS(SEQ_{Agent},SEQ^{*})|}{|SEQ_{Agent}|}$ (1)
where $SEQ_{Agent}$ represents the action sequence generated by the agent, while $SEQ^{*}$ is the ground truth. $LCS(A,B)$ denotes the LCS between $A$ and $B$. $|A|$ is the length of sequence $A$. The keypoint score emphasizes the completeness of the enacted behaviors, ensuring all crucial actions and expressions are performed. The appropriateness score goes beyond completeness to evaluate the overall coherence and naturalness of the performance, considering how well the behaviors flow together and align with the script’s dramatic intent. The human evaluation score is an integer between 1 and 5, with higher scores indicating better performance. It provides a comprehensive assessment of the agents’ ability to not only execute actions accurately but also to deliver them naturally and engagingly. Details of the human evaluation criteria can be found in Section R.
Figure 4: Illustration of the architecture of Alex.
@@ 4 Alex Agent
To truly demonstrate the challenge of this benchmark, we propose Alex [^6] Alex is the protagonist in the sandbox game Minecraft, one of the default skins for players and a character in the game: https://www.minecraft.net/zh-hans. To pay tribute, we named our proposed AI agent Alex., a VLM-based agent as shown in Figure 4. Conventional LLM-powered AI agents depend on LLM to operate as its brain, which is backed by several vital components that perform various essential functions. These components, such as the memory component, planning component, and acting component, have been thoroughly studied recently. To cater to our specific requirements, we have improved these three components (refer to Section E for more details), replaced LLM with VLM and introduced one new component: the multitasking component. Additionally, Alex exhibits different personality traits predefined in the system prompt.
@@
People often switch attention between tasks, for example cooking while talking. The ability to communicate smoothly with other players while working on a task-oriented action is crucial in multi-agent scenarios. Therefore, we develop the mechanism of multitasking ability to enhance the agent’s attention control and working memory abilities inspired by the Multitasking theory from the Cognition field [53]. Specifically, for attention control, the interrupt mechanism effectively controls attention among multiple tasks. For working memory, Alex maintains and processes information in the Memory Library. When another agent says hello to the working agent, this involves saving and restoring internal states when frequent and high-speed switching between communication activities and goal-driven working actions to avoid forgetting ongoing tasks. With the multitasking mechanism, Alex allows for simultaneously simulating and executing intricate coordination and scheduling with multiple tasks.
@@ 5.1 Experiments of Simulators Performance
We evaluate the number of agents that MineLand can support and compare MineLand with other popular Minecraft simulators. We utilize a mainstream consumer desktop PC equipped with an Intel i5-12400F CPU and 64GB of memory. Performance Monitor is employed to monitor the process. Our findings reveal that MineLand is capable of supporting 32 agents simultaneously while providing visual displays. When visual display is disabled, the number of concurrently running agents increases to $\times$4 times. Furthermore, as depicted in Table 3, when MineLand and Malmo both run 8 agents, MineLand’s CPU and memory usage are approximately 1/3 that of Malmo’s (specifically, 35.6% and 38.0%, respectively). It is worth noting that Malmo serves as the foundation for most popular Minecraft Platforms (e.g., MineDojo/MineRL/MarLÖ [42]), thus highlighting MineLand’s superior performance compared to the vast majority of existing Minecraft Platforms. Consequently, MineLand proves to be highly suitable for multi-agent environments.
5.2 Experiments of Social Dynamics
In the “unlocking tools” task with two agents (Table 3), two agents in the collaborative mode work together effectively, with a reduced workload per agent and higher communication expenses. On the other hand, two competitive agents worked independently and necessitated fewer code iterations. The primary reason is that, in competitive relationships, agents tend to achieve more in a single iteration to expedite progress and outperform their opponents. However, this results in less thorough planning and more code errors. Consequently, multiple agents in competitive relationships require fewer code iterations but make more mistakes.
We also observe that personality plays a significant role in determining the behavior of agents in multi-agent societies [25]. We assigned the personality traits of high extraversion and agreeableness to both agents. Under this condition, the agents tended to establish collaboration and engage in mutual communication (co-op >8 out of 10 times). This result is consistent with human behavior [13]. When no personality was set for the agents, they worked independently (co-op 0 out of 10 times). For more experiments of simulating sociological phenomena with >10 agents, please refer to Section O.
[…]
@@ 5.2 Experiments of Social Dynamics
Table 3: The number of code iterations needed per agent to unlock tools made of various materials is determined under three conditions. These conditions include a single agent, two agents in cooperation, and two agents in competition. Each experiment is repeated three times, and the success rate is 100%.
Relationship Material
Wooden Stone Iron
Single Agent 7±2 10±3 25±7
Cooperative 13±5 20±7 49±10
Competitive 6±2 10±3 27±10
5.3 Experiments of Construction Tasks
To establish a baseline for the research community, we evaluate our un-finetuned Alex on two example construction tasks: “Monument Construction” and “Stone Stele Construction” (detailed in Section M and Table 5). While Alex demonstrates a promising capability in selecting appropriate materials for both tasks, its overall construction abilities are currently limited. We posit two key reasons for these limitations: complex task planning and high-precision manipulation. For the first reason, construction tasks are inherently time-consuming and require sophisticated planning abilities. For instance, building auxiliary structures to facilitate the main construction process is crucial for achieving significant height. However, Alex cannot currently plan and construct such auxiliary structures, hindering its ability to construct tall structures. The second reason is the high-precision manipulation. The construction tasks necessitate high-precision manipulation, such as adding decorative details. Alex primarily utilizes high-level actions and lacks the necessary APIs to execute these fine-grained manipulations effectively. These results show the challenge of this task.
5.4 Experiments of Stage Performance Tasks
We evaluate our un-finetuned Alex on four example stage performance tasks (Table 5). We found that in single-agent tasks, the agent typically exhibits high scores across all three evaluation metrics. However, in multi-agent tasks, the keypoint score tends to be higher than the appropriate score because the agent can understand the script and fulfill the key requirements. In contrast, the human evaluation scores are usually relatively lower because agents often engage in unnecessary dialogue in scripts that require collaboration, resulting in lower scores.
Table 4: Human evaluation score and VLM-based (gpt-4o) evaluation score of Alex agent on two construction tasks. We conducted three evaluations using the VLM and obtained consistent results. Details in Section Q. VLM: VLM-based score. Human: Human evaluation score.
Construction Score
VLM Human
A Monument 4, 3, 3 3
A Stone Stele 1, 1, 1 1
Table 5: Appropriateness score (Appro.), keypoint score (Key.), and Human Evaluation Score (Human.) of Alex Agent across four stage performance tasks. The number in parentheses following the script name represents the number of agents in that script. Appropriateness score and keypoint score are real numbers in the range $[0,1]$, and the human evaluation score is an integer between 1 and 5. Details in Section N and Section R.
Script Name Score
Appro. Key. Human.
Cook food(1) 1.00 1.00 5
Exchange items(2) 0.59 0.99 4
Make friends(3) 0.67 0.98 3
Romeo and Julia, Act I Scene I(13) 0.09 0.20 1
5.5 Experiments of Multi-Agent Cooperation
@@ 5.6 Experiments of Multitasking
To validate the impact of multitasking support in the simulator [^7] This multitasking functionality is achieved through the synergistic interplay of two keys: the multitasking component in Alex and the interrupt mechanism in MineLand., we conduct the obsidian mining task, which takes over 8 minutes and requires multiple steps to complete. In the beginning, the agent is mining and encounters chat or hurt events within 8 minutes. Chat event: Another agent nearby initiates a conversation with Alex. This is a low-priority event, and Alex can choose whether to respond to the agent nearby. Hurt event: The agent gets hurts. For example, a zombie attacks the agent. This is a high-priority event, requiring Alex to stop its current task and address this event first.
@@ 6 Conclusion
Traditional multi-agent simulators struggle with large-scale scenarios and often assume perfect information and unrealistic agent capabilities. To address these limitations, we introduce MineLand, a novel Minecraft-based simulator supporting 64 or more agents with limited senses and physical needs. This forces agents to actively communicate and collaborate, fostering more ecologically valid interactions. The advantage carries potential broader impacts across various domains as discussed in Section B.
Appendix A Related Work
A.1 Multi-Agent Simulator
As the popularity of AI agent research continues to grow, there has also been a focus on studying multiple AI agents as well as their cooperation and competition. Researchers and practitioners imagine a dynamic artificial society where human interactions can be simulated by trustworthy agents [12]. From two individuals [9, 45], through four individuals [61], to sandbox games Smallville with twenty-five individuals [40], we witness how individuals perceive a simulated society as the backdrop and interact with the agents and people who engage with it. Each individual can be portrayed through a program, a real human, or an agent based on LLM [40]. The interaction between individuals plays a role in shaping social behavior, leading to simulation of the society. Simulating larger societies can be advantageous. Increasing the number of agents can lead to greater specialization, enabling the accomplishment of more complex and larger-scale tasks. This can significantly improve task efficiency, such as in software development tasks [48]. Additionally, such simulation of interaction has had a significant effect in many other fields. For example, it can replicate realistic social phenomena [14, 41], enhance social robots [4, 6]. They can also be used to test social science theories [5, 24, 23], create model human processors for theory and usability testing [10, 26], train people on how to handle rare yet difficult interpersonal situations [56, 28, 22], and support game characters [32, 52]. There are also many non-Minecraft and non-open-world simulators, such as Habitat3.0 [47], VirtualHome [44, 46], SAPIEN [62], AI2-THOR [31], and iGibson [35, 54].
Challenges of Scaling Up the Number of Agents. While increasing the number of agents can improve task efficiency and make multi-agent simulations more realistic [48, 40, 60], current research primarily focuses on a small number of agents [40, 2, 70]. This is mainly due to the challenges of scaling up the number of agents. Deploying a large number of AI agents will result in an increased computational burden, necessitating better architectural design and computational optimization [40]. Most research in terms of AI agents mimicking daily life routines, focused on two agents [45]. Simulators that remind people of sandbox games (The Sims) initially support four individuals [61], then are extended to twenty-five individuals by [40].
Challenges of Limited Multimodal Senses. To ensure the authenticity of the simulation, an ideal multi-agent simulator should operate under the fundamental assumption that agents possess only limited multimodal senses like humans [21]. Limited multimodal senses mean partially observable environments and an eco-centric perspective. Limited visual and auditory senses restrict information access, forcing agents to actively navigate and communicate to compensate for sensory deficiencies. This mirrors real-life social interactions, where visibility and audibility can be affected by factors such as distance, terrain, and context [29]. As the number of agents grows, the challenges of limited multimodal senses become quite difficult. This is because the communication network of the entire system becomes highly intricate. For agents in our MineLand, the video input is an eco-centric perspective (first perspective) instead of the third perspective in [40], which is omniscient and unrealistic.
Challenges of Physical Needs. Multi-agent simulation platforms hold immense potential for exploring and understanding human social dynamics. However, existing paradigms often disregard the human needs [15, 1]. Existing simulators are designed to simulate believable human behavior in daily-life activities [40, 2]. These activities include waking up, cooking breakfast, heading to work, and initiating conversations with others. However, they do not define their physical needs. For example, after a certain amount of time has passed, the agent will become hungry and have the desire to cook. This desire then leads to the next actions. In this way, the action of cooking is motivated by real desires instead of a predefined schedule.
A.2 Multi-Agent Simulator w.r.t Minecraft
Minecraft, the beloved sandbox game, has been a valuable platform for researchers exploring various fields, including artificial intelligence and multi-agent systems, because of its open world and diverse mechanics.
Specifically, the flexibility and richness of Minecraft make it perfect for developing multi-agent simulators. Researchers have the ability to create various custom environments and scenarios within the game world, where they can introduce virtual agents with specific goals and capabilities. These agents can then interact with each other and the environment, providing researchers with the opportunity to observe and analyze their behavior in a controlled setting. These simulators can be broadly categorized into two main types: task-oriented simulators and daily-life simulators.
Task-oriented simulators focus on agents achieving specific objectives within a set time frame. Minecraft is regarded as the training ground for AI agents to hone their skills. For example, [19, 16, 68, 37] focuses on exploring the environment and gathering resources like wood and stone, and managing them efficiently to complete tasks like building structures or crafting tools. Agents in [67] are designed to complete navigation tasks. In [11], agents are self-organized to execute collaborative building tasks and resource collection tasks. Agents in [18] must work together to overcome challenges that require joint effort.
Daily-life simulators take a more holistic approach, focusing on the daily lives of agents within a virtual society. [40] simulates the rhythms and routines of daily life. This includes activities such as waking up, cooking breakfast, going to work, forming opinions, observing others, and engaging in conversations. [33] simulates sociological experiments, and examines how risk affects the way people engage in bartering.
A.3 AI Agent with LLMs and VLMs
LLMs or VLMs are commonly utilized to bootstrap the components of the Agent. In particular, LLMs have demonstrated effective performance for task-planning [18, 69, 38, 36, 17], and they possess substantial world knowledge [63]. Team CraftJarvis has also developed a powerful agent, [58], based on their previous projects [59, 7]. Moreover, VLMs like CLIP [50] offer a versatile visual-language representation that aligns with language and enables zero-shot visual recognition capabilities for potential AI agents. Using VLMs in construction tasks has many advantages over existing methods [64]. AI agents with visual information can also complete tasks in human-like ways [49, 66].
Appendix B Broader Impact